# On the multiplier spectrum of polynomials

### v1

@@ 1.2. Statement of the main results
Let $d\in\{2,3\}$. There exists a non-empty Zariski open subset $U$ of ${\rm Poly}^{d}(\mathbb{C})$ such that for all $(f,g)\in U\times U$ with $f$ and $g$ intertwined, we have $[f]=[g]$ in ${\rm MPoly}^{d}(\mathbb{C})$.
Proposition 1.6 is also a direct corollary of [FG22, Theorems 3.51 and 3.52] (see also [GNY19, Theorem 1.4]). We include this to present different approaches.
@@
The intermediate results presented in §3 may be useful for studying the dynamics of one-variable polynomials and their products. For instance, we present a proof of the Zariski-dense orbit conjecture for a general split polynomial endomorphism on $(\mathbb{P}^{1})^{2}$ with all factors of degree $d\geq 2$ (see Theorem 3.33).
@@
After establishing the generic injectivity of the multiplier spectrum morphism, a natural further direction is to study the non-injective locus of $\tau_{d}$ (resp. $\tilde{\tau}_{d}$), i.e., the locus
$\{([f],[g])\in\mathcal{M}_{d}(\mathbb{C})^{2}\mid\tau_{d,m_{d}}([f])=\tau_{d,m_{d}}([g])\}\setminus\Delta,$
or its intersection with ${\rm MPoly}^{d}(\mathbb{C})^{2}$, where $\Delta$ is the diagonal in $\mathcal{M}_{d}(\mathbb{C})^{2}$.
There are two known cases of non-conjugate rational maps with the same multiplier spectrum. For a non-constant rational map $f$, if $f=h_{1}\circ h_{2}$ in $\mathbb{C}(z)$, then $h:=h_{2}\circ h_{1}$ is called an elementary transformation of $f$. Two non-constant rational maps $f$ and $g$ are called equivalent, written $f\sim g$, if there exists a finite chain of elementary transformations between $f$ and $g$. Clearly, equivalence defines an equivalence relation for non-constant rational maps, which is coarser than conjugacy (i.e., conjugacy implies equivalence), so we can talk about the equivalence of two conjugacy classes. An easy lemma of Pakovich [Pak19, Lemma 2.1] shows that if $f$ and $g$ are equivalent (of degree $\geq 2$), then they have the same multiplier spectrum. The other known case of non-injectivity comes from exceptional maps. Silverman [Sil07, Theorem 6.62] showed that $\tau_{d}$ is not injective at $[f]$ for some rigid Lattès maps $f$. Pakovich asked [Pak19, Problem 3.1] whether these are the only obstructions to the injectivity of $\tau_{d}$. Ji and Xie [JX25, Conjecture 1.5] conjectured that these are the only obstructions to the injectivity of $\tau_{d}$:
@@
Let $f$ and $g$ be non-conjugate rational maps of degree $d\geq 2$ with the same multiplier spectrum. Then one of the following holds:
@@
When $f$ and $g$ are polynomials, only case ii in Conjecture 1.7 can occur. Note that if the degree $d\geq 2$ is a prime number, then two rational maps $f$ and $g$ of degree $2\leq d\leq 3$ are equivalent if and only if they are conjugate. Consequently, Conjecture 1.7 predicts that for every prime $p\in\mathbb{Q}$, the multiplier spectrum morphism $\tilde{\tau}_{p}$ is injective on ${\rm MPoly}^{p}(\mathbb{C})$. On the other hand, an observation of Huguin [Hug24, Corollary 85] shows that $\tilde{\tau}_{d}$ is not injective for every composite $d\geq 2$.
If the degree $d$ is small, then it is possible to analyze the non-injective locus of $\tilde{\tau}_{d}$ (indeed, $\tilde{\tau}_{d,m}$ with small $m$) using explicit computations. For the prime degree $2\leq d\leq 3$, Huguin [Hug24, Theorem C] showed that $\tilde{\tau}_{d,1}$ is injective on the moduli space ${\rm MPoly}^{d}(\mathbb{C})$. (The injectivity of $\tilde{\tau}_{2,1}$ (in fact, $\tau_{2,1}$) was first proved by Milnor [Mil93].) For the composite degree $d=4$, a result [Hug24, Proposition 86] of Huguin implies that for every $(f,g)\in{\rm Poly}^{4}(\mathbb{C})^{2}$, $\tilde{\tau}_{4}(f)=\tilde{\tau}_{4}(g)$ if and only if $\tilde{\tau}_{4,2}(f)=\tilde{\tau}_{4,2}(g)$ if and only if $[f]=[g]$ in ${\rm MPoly}^{4}(\mathbb{C})$ or $(f,g)=(h_{1}\circ h_{2},h_{2}\circ h_{1})$ for some $(h_{1},h_{2})\in{\rm Poly}^{2}(\mathbb{C})^{2}$ (in particular, $f\sim g$), whose proof relies on computations using a computer.
We intend to investigate the non-injective locus of the multiplier spectrum morphism for polynomials. In §4, for all $d\geq 2$, we provide a first description of the non-injective locus of $\tilde{\tau}_{d}$, i.e., the locus
${\rm PNI}_{d}:=\{([f],[g])\in{\rm MPoly}^{d}(\mathbb{C})^{2}\mid\tau_{d,m_{d}}([f])=\tau_{d,m_{d}}([g]),[f]\neq[g]\},$
which is a starting point of Conjecture 1.7 for the polynomial case:
@@
Let $d\geq 2$ be an integer, and let $C\subset{\rm MPoly}^{d}\times{\rm MPoly}^{d}$ be an irreducible curve defined over $\overline{\mathbb{Q}}$ such that the projection of $C$ to each factor ${\rm MPoly}^{d}$ is non-constant. Assume $C(\mathbb{C})\setminus{\rm PNI}_{d}$ is finite. Then for every $t=([f_{t}],[g_{t}])\in C(\mathbb{C})$, $f_{t}$ and $g_{t}$ are intertwined.
@@
Let $\phi_{1},\phi_{2}:C\times\mathbb{P}^{1}\to C\times\mathbb{P}^{1}$ be two non-isotrivial (algebraic) families over $\overline{\mathbb{Q}}$ of degree-$d$ ($d\geq 2$) polynomials, parameterized by an affine irreducible curve $C$ over $\overline{\mathbb{Q}}$. Denote the induced morphisms also by $\phi_{1},\phi_{2}:C\to{\rm MPoly}^{d}$. Assume $(\phi_{1}(t),\phi_{2}(t))\in{\rm PNI}_{d}$ for all but finitely many $t\in C(\mathbb{C})$. (Here non-isotriviality means that $\phi_{j}(C)$ is not a single point in ${\rm MPoly}^{d}$ for $j=1,2$.)
@@
    there exist $N,k_{1},k_{2},l\in\mathbb{Z}_{>0}$ and $V(z)\in\mathbb{C}[z]\setminus\mathbb{C}$, with k_1k_2,gcd(d,k)=1,l1 (mod k),V(0)0 such that ϕ_1(t)^Nz^lV(z^k_1)^k_1^and ϕ_2(t)^Nz^lV(z^k_2)^k_2^, where $k:={\rm lcm}(k_{1},k_{2})$ denotes the least common multiple of $k_{1}$ and $k_{2}$, and $k_{j}^{\prime}=k/k_{j}$ for $j=1,2$.
@@
$(\phi_{1}(t)^{\circ N})_{\rm min}\sim(\phi_{2}(t)^{\circ N})_{\rm min},$
see Remark 4.4.
In (2), if $k_{j}^{\prime}>1$ ($j=1,2$), then $z^{l}\cdot V(z^{k_{j}})^{k_{j}^{\prime}}$ is a generalized Lattès map in the sense of Pakovich [Pak20c] (see Definition 2.3 and Proposition 2.4). Note that at least one $k_{j}^{\prime}>1$ since $k_{1}\neq k_{2}$.
@@
In Theorem 1.9, we can choose NN in (1) or (2) to have an upper bound depending only on $\mathbb{Q}(\phi_{1},\phi_{2})$ and $d$, see Remark 4.5.
The proof of Theorem 1.9 is divided into two steps. The first step is as follows. After a base change, we may assume that all the critical points are marked for $\phi_{1}$ and $\phi_{2}$. It is well-known that post-critically finite (PCF) parameters do not form families in ${\rm MPoly}^{d}$. We argue that the multiplier spectrum determines the periodicity of critical points; see Proposition 4.3 for a precise statement. Then from $\phi_{1}$ and $\phi_{2}$ we can construct two entangled active dynamical pairs in the sense of Favre-Gauthier [FG22], still denoted by $\phi_{1}$ and $\phi_{2}$ with parameters in $C$. Applying a theorem of Favre-Gauthier (Theorem 4.1), we see that both $\phi_{1}^{\circ N}$ and $\phi_{2}^{\circ N}$ are semi-conjugate to a common family $R$ of polynomials over a non-empty Zariski open subset $U$ of $C$, for some integer $N\geq 1$. In the second step, we need some results of Pakovich on generalized Lattès maps [Pak20c]. Set $S=(C\setminus U)(\mathbb{C})$, which is a finite set. Fix $t\in U(\mathbb{C})$. Assume that (1) does not hold for $t$. Then [Pak20c, Theorem 2.10] implies that $R(t)$ is a generalized Lattès map. By the classification of generalized Lattès maps that are polynomials (Proposition 2.4) and some elementary arguments, we conclude that (2) holds for $t$, after possibly further iteration.
@@ 1.3. Organization of the article
In §2, we provide a brief introduction to orbifolds and generalized Lattès maps based on [Pak20c], for the needs of §3 and §4.
In §3, we explore the generic injectivity of the multiplier spectrum morphism for polynomials. In §3.1, we introduce the notion of pre-simple polynomials and study their properties, essentially done by Pakovich [Pak25a]. Then we deduce Theorem 1.5. In §3.2, we study general polynomials of degree $d\in\{2,3\}$ and obtain Proposition 1.6, following the ideas in [Pak25b]. In §3.3, we deduce Theorem 1.4 from Theorem 1.5 and Proposition 1.6, following Ji-Xie [JX25]. Finally, in §3.4, we present a proof of the Zariski-dense orbit conjecture for a general split polynomial endomorphism on $(\mathbb{P}^{1})^{2}$ with all factors of degree $d\geq 2$ (Theorem 3.33).
In §4.1, we recall the notion of critically marked polynomial and state a theorem of Favre-Gauthier (Theorem 4.1), which is essential for the proof of Theorem 1.9. Then we proceed to complete the proofs of Theorem 1.9 and Theorem 1.8 in §4.2. Subsection 4.3 is devoted to the study of the situation related to Theorem 1.9 (2). We consider pairs of polynomials of the form $(z^{r}\cdot R(z)^{k},z^{r}\cdot R(z^{k}))$, called Ritt moves in this article. We show that under specific conditions $z^{r}\cdot R(z)^{k}$ and $z^{r}\cdot R(z^{k})$ have the same multiplier spectrum over an arithmetic progression (Theorem 4.6) and that they always have the same set of multipliers up to powers (Proposition 4.11).
@@ Acknowledgement
The author would like to express his great gratitude to his advisor, Professor Junyi Xie, for the constant encouragement. The author is supported by NSFC Grant (No. 12271007). The author would like to thank Valentin Huguin for helpful comments on the first version of this paper.
@@ 2. Orbifolds and generalized Lattès maps
This section collects definitions and results concerning orbifolds and generalized Lattès maps, as needed for our purposes. The theory of generalized Lattès maps (based on orbifolds) was primarily developed by Pakovich; see [Pak20c, Pak16, Pak18, Pak20a, Pak20b, Pak23]. Here, we restrict our attention to orbifolds on the compact Riemann surface $\mathbb{P}^{1}(\mathbb{C})$, which suffices for our needs.
@@
An orbifold $\mathcal{O}$ on $\mathbb{P}^{1}(\mathbb{C})=\widehat{\mathbb{C}}$ is defined by a ramification function $\nu:\widehat{\mathbb{C}}\to\mathbb{Z}_{>0}$ such that $\nu^{-1}(\left[2,\infty\right[)$ is discrete. We only consider good orbifolds, i.e., those satisfying $\nu^{-1}(\left[2,\infty\right[)\neq 1$ and, if $\nu^{-1}(\left[2,\infty\right[)=2$, then $\nu(\widehat{\mathbb{C}})=2$. For an orbifold $\mathcal{O}$ on $\widehat{\mathbb{C}}$ with ramification function $\nu$, its Euler characteristic is defined as
@@
the set of singular points is $c(\mathcal{O})=\{z\in\widehat{\mathbb{C}}:\nu(z)>1\}$, and the signature is the multi-set $\nu(\mathcal{O})=\{\nu(z):z\in c(\mathcal{O})\}$ of cardinality $c(\mathcal{O})$.
@@
Clearly, a covering map between orbifolds is also a minimal holomorphic map.
An equivalent description of Lattès maps is well-known: a rational map $A:\mathbb{P}^{1}(\mathbb{C})\to\mathbb{P}^{1}(\mathbb{C})$ of degree $\geq 2$ is a Lattès map if and only if there exists an orbifold $\mathcal{O}$ on $\widehat{\mathbb{C}}$ with a non-constant ramification function such that $A:\mathcal{O}\to\mathcal{O}$ is a covering map between orbifolds. This motivates the following definition:
@@
A rational map $A:\mathbb{P}^{1}(\mathbb{C})\to\mathbb{P}^{1}(\mathbb{C})$ of degree $\geq 2$ is called a generalized Lattès map if there exists an orbifold $\mathcal{O}$ on $\widehat{\mathbb{C}}$ with a non-constant ramification function such that $A:\mathcal{O}\to\mathcal{O}$ is a minimal holomorphic map between orbifolds.
From the definition, it is clear that all exceptional rational maps of degree $\geq 2$ are generalized Lattès maps. From [Pak20c, Proof of Theorem 7.2], we obtain the following complete description of polynomials that are generalized Lattès maps:
@@
    Suppose $f$ is non-exceptional. If $\mathcal{O}$ is an orbifold on $\widehat{\mathbb{C}}$ with a non-constant ramification function such that $f:\mathcal{O}\to\mathcal{O}$ is a minimal holomorphic map between orbifolds, then $\nu(\mathcal{O})=\{n,n\}$ for some integer $n\geq 2$, and there exist a positive integer $r$ with $\gcd(r,n)=1$ and a non-constant polynomial $R\in\mathbb{C}[z]\setminus\mathbb{C}$ such that $f$ is conjugate to $z^{r}\cdot R(z)^{n}$. In fact, $f$ is a generalized Lattès map if and only if there exist integers $r\geq 1$ and $n\geq 2$ with $\gcd(r,n)=1$, and a polynomial $R\in\mathbb{C}[z]$ such that $f$ is conjugate to $z^{r}\cdot R(z)^{n}$.
@@
Let $f,g\in\mathbb{C}(z)\setminus\mathbb{C}$ be rational maps of degree $\geq 2$ such that $f\geq g$. If $f$ is a generalized Lattès map, then so is $g$.
Two natural orbifolds are associated with a rational map:
@@
It is straightforward to verify that a non-constant rational map $f:\widehat{\mathbb{C}}\to\widehat{\mathbb{C}}$ is a covering map $f:\mathcal{O}_{1}^{f}\to\mathcal{O}_{2}^{f}$ between orbifolds.
Let $f:\mathbb{P}^{1}_{\mathbb{C}}\to\mathbb{P}^{1}_{\mathbb{C}}$ be a non-constant rational map over $\mathbb{C}$, i.e., $f\in\mathbb{C}(z)\setminus\mathbb{C}$. A rational map $f_{0}\in\mathbb{C}(z)$ is called a compositional right factor of $f$ if there exists $f_{1}\in\mathbb{C}(z)$ such that $f=f_{1}\circ f_{0}$. Similarly, we define compositional left factors. Pakovich [Pak20c] introduced the notion of good solutions for certain equations, which includes the following special case:
@@
where $f,p,g,q:\mathbb{P}^{1}_{\mathbb{C}}\to\mathbb{P}^{1}_{\mathbb{C}}$ are non-constant rational maps over $\mathbb{C}$. A solution $(f,p,g,q)$ of (2.1) is called good if the fiber product of $f$ and $g$ consists of a unique component, and $p$ and $q$ share no common compositional right factor of degree at least $2$.
Pakovich [Pak20c, Theorem 5.1] described all good solutions of (2.1) with $p(z)=g(z)=z^{n}$ as follows:
@@
Let $n\geq 2$ be an integer and $A,F\in\mathbb{C}(z)$ be rational maps of degree at least $2$ such that $A\circ z^{n}=z^{n}\circ F$. Then the following are equivalent:
@@
    There exist $R\in\mathbb{C}(z)$ and $r\in\mathbb{Z}_{>0}$ with $\gcd(r,n)=1$ such that A(z)=z^rR(z)^n and F(z)=z^rR(z^n).
@@
Pakovich [Pak20c] showed that rational maps that are not generalized Lattès maps exhibit “better” behavior than generalized Lattès maps. The following theorem, a special case of [Pak20c, Theorem 3.2], illustrates this.

### v2

@@ 1.2. Statement of the main results
Let $d\in\{2,3\}$. There exists a non-empty Zariski open subset $U$ of $\mathrm{Poly}^{d}(\mathbb{C})$ such that for all $(f,g)\in U\times U$, $f$ and $g$ are intertwined if and only if $[f]=[g]$ in $\mathrm{MPoly}^{d}(\mathbb{C})$.
Proposition 1.7 is also a direct corollary of [FG22, Theorems 3.51 and 3.52] (see also [GNY19, Theorem 1.4]). We include this proposition to present different approaches.
@@
The intermediate results presented in § 3 may be useful for studying the dynamics of one-variable polynomials and their products. For instance, we present a proof of the Zariski-dense orbit conjecture for a general split polynomial endomorphism on $\mathbb{A}^{2}$ with all factors of degree $d\geq 2$ (see Theorem 3.39).
@@
After establishing the generic injectivity of the multiplier spectrum morphism, a natural further direction is to study the non-injective locus of $\tau_{d}$ (or $\tilde{\tau}_{d}$), i.e., the locus
$\{([f],[g])\in\mathcal{M}_{d}(\mathbb{C})^{2}\mid\tau_{d,m_{d}}(f)=\tau_{d,m_{d}}(g)\}\setminus\Delta,$
or its intersection with $\mathrm{MPoly}^{d}(\mathbb{C})^{2}$, where $\Delta$ is the diagonal in $\mathcal{M}_{d}(\mathbb{C})^{2}$.
Outside the flexible Lattès locus, there are two known cases of non-conjugate isospectral rational maps. For a non-constant rational map $f\in\mathbb{C}(z)\setminus\mathbb{C}$, if $f=h_{1}\circ h_{2}$ for some $h_{1},h_{2}\in\mathbb{C}(z)$, then $h:=h_{2}\circ h_{1}$ is called an elementary transformation of $f$. Two non-constant rational maps $f$ and $g$ are called equivalent, written $f\sim g$, if there exists a finite chain of elementary transformations between $f$ and $g$. Clearly, equivalence defines an equivalence relation on $\mathbb{C}(z)\setminus\mathbb{C}$, which is coarser than conjugacy (i.e., conjugacy implies equivalence). So we can talk about the equivalence of two conjugacy classes of rational maps. An easy lemma of Pakovich [Pak19b, Lemma 2.1] shows that if two rational maps $f$ and $g$ are equivalent (of degree $\geq 2$), then they are isospectral. The other known case of non-injectivity comes from exceptional maps. Silverman [Sil07, Theorem 6.62] showed that $\tau_{d}$ is not injective at $f]\in\mathcal{M}_{d}(\mathbb{C})$ for some rigid Lattès $f$. Pakovich asked [Pak19a, Problem 3.1] whether these are the only obstructions to the injectivity of $\tau_{d}$. Ji and Xie [JX25, Conjecture 1.5] conjectured that Pakovich’s question has a positive answer:
@@
Let $f$ and $g$ be non-conjugate isospectral rational maps of degree $d\geq 2$. Then one of the following holds:
@@
Since polynomials are not Lattès, we only need to consider the case (ii) in Conjecture 1.8. Note that if the degree $d$ is a prime number, then two rational maps $f$ and $g$ of degree $d$ are equivalent if and only if they are conjugate. Consequently, Conjecture 1.8 predicts that for every prime $p$, the multiplier spectrum morphism $\tilde{\tau}_{p}$ is injective on $\mathrm{MPoly}^{p}(\mathbb{C})$. On the other hand, an observation of Huguin [Hug24, Corollary 85] shows that $\tilde{\tau}_{d}$ is not injective on $\mathrm{MPoly}^{d}(\mathbb{C})$ for every composite $d\geq 2$.
If the degree $d$ is small, then it is possible to analyze the non-injective locus of $\tilde{\tau}_{d}$ (indeed, $\tilde{\tau}_{d,m}$ with small $m$) using explicit computations. For prime degree $2\leq d\leq 3$, Huguin [Hug24, Theorem C] showed that $\tilde{\tau}_{d,1}$ is injective on the moduli space $\mathrm{MPoly}^{d}(\mathbb{C})$. Recall that Milnor [Mil93] showed that $\tau_{2,1}$ is injective on $\mathrm{Rat}_{2}(\mathbb{C})$. For the composite degree $d=4$, a result [Hug24, Proposition 86] of Huguin (whose proof relies on computer computations) shows that the following are equivalent for all $(f,g)\in\mathrm{Poly}^{4}(\mathbb{C})^{2}$:
  • $\tilde{\tau}_{4}(f)=\tilde{\tau}_{4}(g)$;
  • $\tilde{\tau}_{4,2}(f)=\tilde{\tau}_{4,2}(g)$;
  • $[f]=[g]$ in $\mathrm{MPoly}^{4}(\mathbb{C})$ or $(f,g)=(h_{1}\circ h_{2},h_{2}\circ h_{1})$ for some $(h_{1},h_{2})\in\mathrm{Poly}^{2}(\mathbb{C})^{2}$ (in particular, $f\sim g$).
We investigate the non-injective locus of the multiplier spectrum morphism $\tilde{\tau}_{d}$ for polynomials. In § 4, for all $d\geq 2$, we provide a first description of the non-injective locus of $\tilde{\tau}_{d}$, i.e., the locus
$\mathrm{PNI}_{d}:=\{([f],[g])\in\mathrm{MPoly}^{d}(\mathbb{C})^{2}\mid\tilde{\tau}_{d}([f])=\tilde{\tau}_{d}([g]),[f]\neq[g]\},$
which serves as a starting point for Conjecture 1.8 for the polynomial case:
@@
Let $d\geq 2$ be an integer, and let $C\subset\mathrm{MPoly}^{d}\times\mathrm{MPoly}^{d}$ be an irreducible curve defined over $\overline{\mathbb{Q}}$ such that the projection of $C$ to each factor $\mathrm{MPoly}^{d}$ is non-constant. Assume that $C(\mathbb{C})\setminus\mathrm{PNI}_{d}$ is finite. Then for every $t=([f_{t}],[g_{t}])\in C(\mathbb{C})$, $f_{t}$ and $g_{t}$ are intertwined.
@@
Let $\phi_{1},\phi_{2}:C\times\mathbb{P}^{1}\to C\times\mathbb{P}^{1}$ be two non-isotrivial (algebraic) families over $\overline{\mathbb{Q}}$ of degree $d$ ($d\geq 2$) polynomials, parametrized by an affine irreducible curve $C$ over $\overline{\mathbb{Q}}$. Denote the induced morphisms also by $\phi_{1},\phi_{2}:C\to\mathrm{Poly}^{d}$. Assume that $([\phi_{1}(t)],[\phi_{2}(t)])\in\mathrm{PNI}_{d}$ for all but finitely many $t\in C(\mathbb{C})$. (Here, non-isotriviality means that $\phi_{j}(C)]$ is not a single point in $\mathrm{MPoly}^{d}$ for $1\leq j\leq 2$.)
@@
    there exist $N,k_{1},k_{2},l\in\mathbb{Z}_{>0}$ and $V(z)\in\mathbb{C}[z]\setminus\mathbb{C}$, with k_1k_2,gcd(d,k)=1,l1 (mod k),V(0)0 such that ϕ_1(t)^Nz^lV(z^k_1)^k_1^and ϕ_2(t)^Nz^lV(z^k_2)^k_2^, where $k:=\mathrm{lcm}(k_{1},k_{2})$ denotes the least common multiple of $k_{1}$ and $k_{2}$, and $k_{j}^{\prime}=k/k_{j}$ for $1\leq j\leq 2$.
@@
$(\phi_{1}(t)^{\circ N})_{\mathrm{min}}\sim(\phi_{2}(t)^{\circ N})_{\mathrm{min}};$
see Remark 4.4 for the precise meaning.
In (2), if $k_{j}^{\prime}>1$ ($1\leq j\leq 2$), then $z^{l}\cdot V(z^{k_{j}})^{k_{j}^{\prime}}$ is a generalized Lattès map in the sense of Pakovich [Pak20c] (see Definition 2.3 and Proposition 2.4). Note that at least one $k_{j}^{\prime}>1$ because $k_{1}\neq k_{2}$.
@@
Let $K$ be a number field such that $\phi_{1}$, $\phi_{2}$, and $C$ are defined over $K$. Then in Theorem 1.10 (1) or (2), we can choose NN to be bounded by a constant depending only on $K:\mathbb{Q}]$ and $d$; see Remark 4.5.
The proof of Theorem 1.10 is divided into two steps. The first step is as follows. After a base change, we may assume that all the critical points of $\phi_{1}$ and $\phi_{2}$ are marked. It is well-known that postcritically finite (PCF) parameters do not form families in $\mathrm{MPoly}^{d}$. We can argue that the multiplier spectrum determines the periodicity of critical points; see Proposition 4.3 for a precise statement. Then from $\phi_{1}$ and $\phi_{2}$ we can construct two entangled active dynamical pairs in the sense of Favre and Gauthier [FG22], still denoted by $\phi_{1}$ and $\phi_{2}$ parametrized by $C$. Applying a theorem of Favre and Gauthier (Theorem 4.1), there exists a family $R$ of polynomials parametrized by a non-empty Zariski open subset $U\subseteq C$ that is semi-conjugate to $\phi_{1}^{\circ N}$ and $\phi_{2}^{\circ N}$, for some integer $N\geq 1$. In the second step, we need some results of Pakovich on generalized Lattès maps [Pak20c]. Set $S=(C\setminus U)(\mathbb{C})$, which is a finite set. Fix $t\in U(\mathbb{C})$. Assume that (1) does not hold for $t$. Then a result of Pakovich [Pak20c] (see Theorem 2.10) implies that $R(t)$ is a generalized Lattès map. By the classification of generalized Lattès maps that are polynomials (Proposition 2.4) and some elementary arguments, we conclude that (2) holds for $t$, after possibly further iteration.
@@ 1.3. Organization of the article
In § 2, we provide a brief introduction to orbifolds and generalized Lattès maps based on [Pak20c], as needed for § 34.
In § 3, we explore the generic injectivity of the multiplier spectrum morphism for polynomials. In § 3.1, we introduce the notion of pre-simple polynomials and study their properties, which are essentially done by Pakovich [Pak25a]. Then we deduce Theorem 1.5. In § 3.2, we study general polynomials of degree $2\leq d\leq 3$ and obtain Proposition 1.7, following the ideas in [Pak25b]. In § 3.3, we deduce Theorem 1.4 from Theorem 1.5 and Proposition 1.7, following Ji and Xie [JX25]. Finally, in § 3.4, we present a proof of the Zariski-dense orbit conjecture for a general split polynomial endomorphism on $\mathbb{A}^{2}$ with all factors of degree $d\geq 2$ (Theorem 3.39).
In § 4.1, we recall the notion of critically marked polynomials and state a theorem of Favre and Gauthier (Theorem 4.1), which is essential for the proof of Theorem 1.10. Then we proceed to complete the proofs of Theorems 1.10 and 1.9 in § 4.2. We study polynomial pairs related to Theorem 1.10 (2) in § 4.3. Precisely, we consider polynomial pairs of the form $(z^{r}R(z)^{k},z^{r}R(z^{k}))$, called Ritt moves in this article. We show that $z^{r}R(z)^{k}$ and $z^{r}R(z^{k})$ have the same multiplier spectrum over an arithmetic progression under specific conditions (Theorem 4.6) and that they always have the same set of multipliers up to powers (Proposition 4.11).
@@ Acknowledgement
The author would like to express his great gratitude to his advisor, Professor Junyi Xie, for the constant encouragement. The author is supported by NSFC Grant (No. 12271007). The author would like to thank Fedor Pakovich and Valentin Huguin for helpful comments on the first version of this paper.
@@ 2. Orbifolds and generalized Lattès maps
This section collects definitions and results concerning orbifolds and generalized Lattès maps necessary for our purposes. The theory of generalized Lattès maps (based on orbifolds) was primarily developed by Pakovich; see [Pak20c, Pak16, Pak18, Pak20a, Pak20b, Pak23]. Here, we restrict our attention to orbifolds on the compact Riemann surface $\widehat{\mathbb{C}}=\mathbb{P}^{1}(\mathbb{C})$, which suffices for our purposes.
@@
An orbifold $\mathcal{O}$ on $\widehat{\mathbb{C}}$ is defined by a ramification function $\nu:\widehat{\mathbb{C}}\to\mathbb{Z}_{>0}$ such that $\{z\in\widehat{\mathbb{C}}\colon\nu(z)>1\}$ is finite. We consider only good orbifolds, i.e., those satisfying $\{z\in\widehat{\mathbb{C}}\colon\nu(z)>1\}\neq 1$, and if $\{z\in\widehat{\mathbb{C}}\colon\nu(z)>1\}=\{z_{1},z_{2}\}$ has cardinality $2$, then $\nu(z_{1})=\nu(z_{2})$. For an orbifold $\mathcal{O}$ on $\widehat{\mathbb{C}}$ with ramification function $\nu$, its Euler characteristic is defined as
@@
the set of singular points is $c(\mathcal{O})=\{z\in\widehat{\mathbb{C}}\colon\nu(z)>1\}$, and the signature is the multiset $\nu(\mathcal{O})=\{\nu(z)\colon z\in c(\mathcal{O})\}$ of cardinality $c(\mathcal{O})$. The orbifold $\mathcal{O}$ is called ramified if its ramification function $\nu$ is not the constant function $1$ on $\widehat{\mathbb{C}}$.
@@
Clearly, every covering map between orbifolds is also a minimal holomorphic map.
A well-known equivalent description of Lattès maps is that a rational map $A:\widehat{\mathbb{C}}\to\widehat{\mathbb{C}}$ of degree $\geq 2$ is a Lattès map if and only if there exists a ramified orbifold $\mathcal{O}$ on $\widehat{\mathbb{C}}$ such that $A:\mathcal{O}\to\mathcal{O}$ is a covering map between orbifolds. This motivates the following definition:
@@
A rational map $A:\widehat{\mathbb{C}}\to\widehat{\mathbb{C}}$ of degree $\geq 2$ is called a generalized Lattès map if there exists a ramified orbifold $\mathcal{O}$ on $\widehat{\mathbb{C}}$ such that $A:\mathcal{O}\to\mathcal{O}$ is a minimal holomorphic map between orbifolds.
From the definition, it is easy to check that all exceptional rational maps of degree $\geq 2$ are generalized Lattès maps. Based on the proof of [Pak20c, Theorem 7.2], we obtain the following complete description of polynomials that are generalized Lattès maps:
@@
    Suppose $f$ is non-exceptional. If $\mathcal{O}$ is a ramified orbifold on $\widehat{\mathbb{C}}$ such that $f:\mathcal{O}\to\mathcal{O}$ is a minimal holomorphic map between orbifolds, then $\nu(\mathcal{O})=\{n,n\}$ for some integer $n\geq 2$, and there exist $r\in\mathbb{Z}_{>0}$ with $\gcd(r,n)=1$ and a polynomial $R\in\mathbb{C}[z]\setminus\mathbb{C}$ such that $f$ is conjugate to $z^{r}R(z)^{n}$. In fact, $f$ is a generalized Lattès map if and only if there exist integers $r\geq 1$ and $n\geq 2$ with $\gcd(r,n)=1$, and a polynomial $R\in\mathbb{C}[z]$ such that $f$ is conjugate to $z^{r}R(z)^{n}$.
@@
Let $f,g\in\mathbb{C}(z)\setminus\mathbb{C}$ be rational maps of degree $d\geq 2$ such that $f\geq g$. If $f$ is a generalized Lattès map, then so is $g$.
One can associate two natural orbifolds with every rational map:
@@
For any non-constant rational map $f:\widehat{\mathbb{C}}\to\widehat{\mathbb{C}}$, the map $f:\mathcal{O}_{1}^{f}\to\mathcal{O}_{2}^{f}$ is clearly a covering map between orbifolds.
Let $f\in\mathbb{C}(z)\setminus\mathbb{C}$ be a non-constant rational map. A rational map $f_{0}\in\mathbb{C}(z)$ is called a compositional right factor of $f$ if there exists $f_{1}\in\mathbb{C}(z)$ such that $f=f_{1}\circ f_{0}$. Similarly, we define compositional left factors. Pakovich [Pak20c, § 2] introduced the notion of good solutions for certain equations:
@@
where $f,p,g,q\in\mathbb{C}(z)\setminus\mathbb{C}$. A solution $(f,p,g,q)$ of (2.1) is called good if $p$ and $q$ share no common compositional right factor of degree at least $2$, and the fiber product of $f$ and $g$ consists of a unique component.
Pakovich [Pak20c, Theorem 5.1] described good solutions of (2.1) with $p(z)=g(z)=z^{n}$ ($n\geq 2$) as follows:
@@
Let $n\geq 2$ be an integer and let $A,F\in\mathbb{C}(z)$ be rational maps of degree at least $2$ such that $A\circ z^{n}=z^{n}\circ F$. Then the following are equivalent:
@@
    There exist $R\in\mathbb{C}(z)\setminus\mathbb{C}$ and $r\in\mathbb{Z}_{>0}$ with $\gcd(r,n)=1$ such that A(z)=z^r R(z)^n and F(z)=z^r R(z^n).
@@
Pakovich [Pak20c] showed that rational maps that are not generalized Lattès maps exhibit “better” behavior than generalized Lattès maps. The following theorem, a special case of [Pak20c, Theorem 3.2], illustrates this: